Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Help!!!

Expert replies
by anuu » Thu Apr 07, 2011 9:04 am
Probability of an employee to get infected by a disease X is 20%. What is the probability that out of randomly selected four employees, 2 employees will suffer from disease X?


Not sure whats the answer for this problem. But here's my approach:

Out of 4 ,2 will suffer from disesse x implies only (exactly) 2:

so lets say X & y are the probabilities an employee 'suffering' and 'not suffering' from disease resp:

therefore x = 20/100 = 1/5 and y = 80/100 = 4/5
so we can have XXYY (the employees can be arranged in 4!/2/!2! ways = 6ways)

(1/5*1/5 *4/5*4/5) = 16/625
Since there r 6 ways total prob = 16/625 + 16/625...6times

i.e 96/625 = 0.1536
Is my approach correct?

I came across this problem in website "urch". there was no solotion posted, so i wanted to check if my approach is rite. Pls correct me if i'm wrong.

Regards,
Anu
Join the discussion
Source: — Problem Solving |

by gmat-challenger » Thu Apr 07, 2011 10:18 am
I think you are perfectly right. that's what I got as an answer.

4!/2!*2! * 1/5^2*4/5*2 = 6*1/25*16/25=96/625

nice problem,
take care
Join the discussion

by rohu27 » Thu Apr 07, 2011 10:40 am
why are we calculating the probability for ONLY 2 employees to suffer and assuming the rest do not suffer?
cant it be that 2 WILL suffer implies, 3 may also suffer or all the four?
2 WILL suffer =2 WILL NOT suffer? i dnt get this :(
Join the discussion

by anuu » Thu Apr 07, 2011 11:05 am
The problem states what is the probability that out of 4 employees ,2 will suffer from infection..it coudl be any 2 out of the four.

Your solution wud worked if the problem stated "ATLEAST 2"..

hope this clears your point..

Anuu
Join the discussion

by force5 » Thu Apr 07, 2011 11:40 am
Yes you almost got it correct. Here is my approach..
this concept is called conditional probability.

Here we are selecting 2 employees and the other two will automatically get selected. HOW?

lets start..

there are 4 employees. now we need to select 2 out of 4.
part-1..... how to select 2 out of 4... simple 4C2 = 6

part 2..... how will 2 employees be selected out of 4.

SS
SS (reverse)
SN
NS
NN
NN (reverse)
(S= suffer , N=Not suffer)

Can there be any other combination..... ??.........NO

Now when you pick SS then NN automatically gets picked. (WHY?? because you can only select 2 employees who will suffer from disease as per question)
similarly when you pick SN then ..... NS /SN
when..........SS then .....NN

now probability of Suffer is 1/5 and probability of not suffer is 4/5
hence in each case irrespective of the position of N or S the total probability will be
1/5 * 1/5 * 4/5 * 4/5= 16/625
hence total probability will be 6* 16/625 = 96/625

some books also call it conditional probability of events occurring p times in t sequence..

tCp * (PROB OF EVENT A)^P * (PROB OF EVENT B)^t-p

AGAIN I WOULD RECOMMEND THE LONGER WAY FOR THE STARTERS. you can derive this result from the above solution too.

Hope it helps...
Join the discussion